OpenStudy (anonymous):

solve the quadratic equation (2x+1)(x-3)=0

OpenStudy (anonymous):

x= -1/2 U x= 3

OpenStudy (anonymous):

Step by step solution : Step 1 : Multiply 2x+1 by x-3 Equation at the end of step 1 : (2x + 1) • (x - 3) = 0 Step 2 : Solve (2x+1)•(x-3) = 0 A product of several terms equals zero. When a product of two or more terms equals zero, then at least one of the terms must be zero.We shall now solve each term = 0 separately. Solving a Single Variable Equation : Solve : 2x+1 = 0 Subtract 1 from both sides of the equation : 2x = -1 Divide both sides of the equation by 2: x = -1/2 Solving a Single Variable Equation : Solve : x-3 = 0 Add 3 to both sides of the equation : x = 3 Two solutions : x = 3 x = -1/2

OpenStudy (anonymous):

@CooperP does this help? u get the answer now?

OpenStudy (anonymous):

Yeah Thanks a lot